# OnlyFlow: Optical Flow based Motion Conditioning for Video Diffusion Models

### v1

@@ 4.2 Model training and inference
We train the model for approximately 20 hours using 8 A100 GPUs. The training process involved a batch size of $32$, a constant learning rate of $1\times 10^{-4}$ with the Adam optimizer. The model consists of 198M parameters in the optical flow encoder $\Phi$ and 19M for the injection layers in the U-Net attention in OnlyFlow-T (resp. 31M for OnlyFlow-ST variant).
Using the trained model
During inference, we can use our model by sample noise in the latent with a text prompt and an auxiliary video. It generates a video whose content best represents the content of the text prompt, while following the dynamics of the motion of the auxiliary video. The two input conditions interact via Eq. (2), where the $\gamma$ hyper-parameter manages the influence of each. Our training was carried out for $\gamma=1.0$, so this is the nominal value for using our model in inference. However, it is possible to vary this parameter between 0 (the optical flow is not taken into account) and 1 to play with the progressive inclusion of motion. This could also enable us to investigate how our model works, by offering comparative videos according to $\gamma$ strength.
With our framework, we can process videos of up to 16 frames. However, there are a number of recent proposals [27, 35], to generate an arbitrary long video without any additional training costs.
5 Experimental Results
5.1 Quantitative results
@@
As a quick recap, our main goal is to transfer the motion cues from a video $\mathbf{V}$ to the generated one $\mathbf{\hat{V}}$ while remaining realistic and faithful to the textual prompt. Therefore, we assess our model quantitatively using three main components: realism, flow fidelity, and textual alignment.
First, to ensure that our model produces realistic videos, we report the mean Fréchet Video Distance (FVD) [34] across frames. This metric evaluates the temporal coherence of generated videos compared to real ones from a distributional perspective. The FVD between a set of real and generated videos, $R$ and $G$, respectively, is given by
$\begin{split}FV&D(R,G)=\\ &|\mu_{R}-\mu_{G}|+\text{Tr}\left(\Sigma_{R}+\Sigma_{G}-2(\Sigma_{R}\Sigma_{G})^{\frac{1}{2}}\right),\end{split}$ (4)
where $\mu_{I}$ and $\Sigma_{I}$ are the mean and covariance of the video encodings in set $I\in{R,G}$. We follow standard practices proposed by Unterthiner et al. [34] to compute the FVD.
Next, we assess the flow fidelity between the generated video and the prompted one, i.e. whether the produced video contains the same motion patterns as the input. To evaluate this, we measure the absolute pixel-wise difference between the optical flow of the input video and the generated data. The error measurement is given by
$\mathbb{E}_{\mathbf{V},F,H,W}\left[\left|\mathcal{T}(\mathbf{V})-\mathcal{T}(\hat{\mathbf{V}})\right|\right],$ (5)
where $\mathcal{T}$ is the optical flow extractor, RAFT in this case.
Finally, to assess the textual alignment between the input text and the generated video, we adopt the CLIP similarity (CS). Specifically, we focus on the expected CS across frames, as described in the next equation:
$\text{CS}\left(V,p\right)=\mathbb{E}_{F}\left[max\left(100\cos\left(E_{\hat{\mathbf{V}}},E_{t}\right),0\right)\right],$ (6)
where $E_{t}$ is the textual CLIP embedding, while $E_{\hat{\mathbf{V}}}$ is the CLIP vision embedding of the generated video. To study the CS, we used a OpenAI Vision Transformer (ViT) architecture as the image encoder, with a patch size of 16x16 pixels [28].
Empirical Results
To empirically validate OnlyFlow, we perform extensive testing using a subset of 70,000 unseen videos from the WebVid [1] dataset to condition our proposed model. The result are shown in Fig. 4. For a fair comparison with standard T2V generative models, we generate videos with $F=16$ frames at a resolution of $512\times 512$ pixels. In addition, to analyze the impact of our proposed motion module, we evaluate different optical flow conditioning strengths $\gamma$ (Eq. (2)).
First, we discuss the FVD metric in Fig.4(a). Interestingly, we notice that as we increase optical flow conditioning strength $\gamma$, the FVD decreases. We posit that this is explained by the addition of real data as a condition to video generation. Real data brings realistic motion which will end up in being used by OnlyFlow, which in turn will be interpreted as accurate by the FVD metric. In this generative process, enforcing the optical flow of an input video with realistic motion thus naturally improves FVD.
@@
Finally, we examine the textual alignment results in Fig. 4(c). The outcome of this experiment show that the model produces stable CLIP scores when using any level of optical flow conditioning. This suggests that the content generated by the model is aligned with the text to the same extent, whether the motion cues are used or not.
5.2 Qualitative results
We now assess OnlyFlow from a qualitative perspective and compare them with previous works. To do so, we present different generative results using the OnlyFlow-T version.
Optical flow conditioning strength
Figure 5: Illustration of OnlyFlow’s optical flow conditioning strength impact on generated videos. Setting the $\gamma$ scale at which the features from the input video are injected sets the influence of its motion on the generated video. The videos presented in each row are obtained for increasing values of $\gamma$ between 00 and $1.0$. We observe a progressive alignment of the generated video motion and wind turbine position on the auxiliary video
Figure 6: Qualitative comparison of video-to-video generation models. Videos are generated using same text prompt and input video. OnlyFlow exhibits a superior combination of motion fidelity and image realism. It positively compares to approaches that uses depth map (RAVE, VideoComposer, Control-A-Video), is comparable to Gen-1 [12]’s temporal coherence and VideoComposer [37]’s image quality.
Comparison with state-of-the-art video-to-video generation models
For the loosely-defined problem of video-to-video editing and customization, we show that our OnlyFlow model performs better for tasks that require only the motion of the input video. The visual comparison in Fig. 6 reveals that OnlyFlow can generate animations whose movements are close to the input video, given a different textual prompt. In this comparison, the rippling waves effect of the auxiliary video struggle to be accurately represented in other models, except for Control-A-Video and Gen-1. We notice that the prompt adherence is also better followed in our approach, Gen-1, VideoComposer and Tune-A-Video. Nevertheless, the latter two models did not achieve to take motion inspiration from our auxiliary video and Gen-1 seems to stack mountains and ripples visuals. Compared to other work such as Tune-A-Video [39], our approach is lightweight. On top of that, it does not require any costly DDIM [32] inversion process used in TokenFlow [13] and RAVE [23].
Figure 7: Qualitative comparison on camera-movement controlled video generation. All videos are generated using the same text prompt, camera trajectory (pan-left), and seed when available. Without having trained for this task, our OnlyFlow achieve the same camera control capability as other camera-movement approaches trained on this task.
In recent work, such as MotionCtrl [38] and CameraCtrl [17], the conditioning of camera motion is achieved with a camera position encoding, while the AnimateDiff [14] motion conditioning uses a LoRA trained on the temporal layer for a given motion. For camera displacements, OnlyFlow provides a flexible alternative for modeling camera movement. To do this, we create an optical flow to describe the desired motion or provide a video with the desired motion. In Fig. 7 we show a specific scenario where we prompted our model with a constant horizontal movement to the left. The generated results clearly show that the generation is on par with the controllability of specific approaches such as CameraCtrl or MotionCtrl. Note that we did not observe satisfactory results with AnimateDiff’s motion LoRA, both in terms of aesthetics and expected camera movements. This shows the versatility of our model.
Semantic alignment
We noticed an interesting capability of semantic alignment between the descriptive information contained in the optical flow fed into the model and the given prompt. As shown in Fig. 8, the optical flow inadvertently captures object shapes. When there is a match between the input video motion concept and the prompt, the model successfully maps the zones of interest to be generated. This ability is similar to other conditioning models such as ControlNet [43] and T2I-Adapter [25], but our method does not require manual control map creation.
(a) Frames from the auxiliary video of which optical flow is extracted
(b) Frames from corresponding generated video
Figure 8: Example of semantic alignment between optical flow and prompt. With the prompt ”Trees in forest”, our OnlyFlow model generates a clothesline corresponding to the teeth of the smile.
User preference study
We employ the Two-alternative Forced Choice (2AFC) protocol for text-driven video editing [2, 11, 13, 26, 40]. Participants are presented with the input video, our results and a baseline, and are asked to determine which video better aligns with the text prompt, which video better preserves the motion of the input video, and which video they prefer overall. We collected 300 user judgments. As seen in Tab.1, our method is consistently preferred over all baselines.
@@ 5.2 Qualitative results
color=yellow!60,inline]H:
Parler des results des 2 types de train
Quantitatif: - diff optical flow - FVD - Clip Score
- CLIP similarity among consecutive video frames (optionnel)
- courbe de training (optionnel) ou annexe
Figure belle pour montrer le comportement de OnlyFlow et que ça marche bien
- flow strength figure image ¿¿ exemple random de comment OnlyFlow marche (avec un exemple juste joli, rien de spécial aka exemple canonique (a remettre dans la figure d’intro du papier)) (2/3 frames + 5 valeurs strength + 2/3 vidéos)
Figure fonctionnalités out of the box
Comparaisons qualitatites avec autres algos:
- CameraCtrl ex avoir des pan ou certains type de mouvements (on compare cet algo avec le fait de conditionner OnlyFlow par des videos de pan et on montre que OnlyFlow sait faire du cameractrl, sans pouvoir avoir été entrainé que à faire ça (point positif))
(comp algo MotionCtrl, CameraCtrl, AD LORA)
- ControlNet ex avoir idem mais ce qu’il peut faire ou non (on compare cet algo avec le fait de conditionner OnlyFlow par des vidéos de type mask que l’on donnerait à controlnet, et on montre que OnlyFlow sait faire du ctrlnet, sans pour autant avoir été entrainé à en faire).
Figure pour montrer que OnlyFlow s’interface bien avec les autres types de conditionnement
[[[[ sparse ctrl, ipadapter, lora ]]]]
Montrer: - ok style transfer - ok pour tache type ControlNet - contrôle de camera
color=teal!60,inline]M: Messages principaux figures resultats: - figure clipscore : ca part pas en vrille: ca ne deteriore pas le suivi du prompt - diff optical flow cut à 0 et baisse - FVD : ca baisse un peu: donc plus fidele, amelioration
6 Discussion and limitations
color=teal!60,inline]M: - lightweight - plus pour gratuits par rapport à - generalisation de controlnet et camera control dans un framework unique (meme si tâche specifique, ex : cameractrl, MotionCtrl)
Discussion
@@
Another motivation for OnlyFlow is the creation of artistic videos. Using the natural motion present in video footage allows the generation of videos that would require tremendous editing efforts to be created without OnlyFlow. In the supplementary video, we provide examples of artistic videos we created during the project, that incorporate upscaling with ESRGAN [36] and frame interpolation with FILM [29], which by themselves showcase the type of outputs that can be done using OnlyFlow. Moreover, OnlyFlow is readily compatible with extensions of AnimateDiff and Stable-Diffusion that are often used in artistic creation context such as IP-Adapter [41], which can further be used to improve the quality and controllability of the generated videos.
@@
One clear limitation of our work is photorealism obtained in the generated videos. While T2V models like AnimateDiff [14] are powerful generative models, their generation remain poor for most complex prompts depicting scenes with various objects and subjects or scenes with movements that might not appear in the training set. The resolution of the generated video are also limited by the underlying model.
@@
7 Conclusion
In this paper, we introduced our OnlyFlow method, which improves on T2V models such as AnimateDiff by conditioning it on the motion extracted from an input auxiliary video. The optical flow of the auxiliary video is then passed to a trainable flow encoder connected to the T2V model. The flow encoder outputs features maps which are injected into the attention layers of the T2V’s U-Net. The generated video is encouraged to accurately follow the motion of the input video. While other methods allow conditioning by movement (with motion masks or user-defined vectorial camera movements), OnlyFlow allows to generate any video with a text prompt and the motion of an auxiliary video. In our experiments, we demonstrate the efficiency of OnlyFlow for various applications such as camera motion control (panning, zooming, etc.) and video editing/style transfer. Without necessarily being trained for such applications, OnlyFlow compares positively to other state-of-the-art methods that have been specifically trained for such applications. We also carried out a user study which confirmed the interest of our method in generating very interesting videos, demonstrating that their movement convincingly and coherently follows the input reference.

### v2

@@ 4.1 Implementation details
We train OnlyFlow for approximately 20 hours using 8 A100 GPUs. The training process involved a batch size of $32$, a constant learning rate of $1\times 10^{-4}$ with the Adam optimizer. The model consists of 198M parameters in the optical flow encoder $\Phi$ and 19M for the injection layers in the U-net attention. To extract the optical flow $f$ during the training phase, we employed the commonly used RAFT-Large [39] model due to its fast inference and widely usage in computer vision tasks.
@@
As a quick recap, our main goal is to transfer the motion cues from a video to the generated one while remaining realistic and faithful to the textual prompt. Therefore, we assess our model quantitatively using three main components: realism, flow fidelity, and textual alignment.
First, to ensure that our model produces realistic videos, we report the mean Fréchet Video Distance (FVD) [40] across frames. This metric evaluates the temporal coherence of generated videos compared to real ones from a distributional perspective. We follow standard practices proposed by Unterthiner et al[40] to compute the FVD.
Next, we assess the flow fidelity between the generated video and the prompted one, i.e. whether the produced video contains the same motion patterns as the input. To evaluate this, we measure the absolute pixel-wise difference between the optical flow of the input video and the generated data.
Finally, to assess the textual alignment between the input text and the generated video, we adopt the CLIP similarity (CS). Specifically, we focus on the expected CS across frames. To study the CS, we used a OpenAI Vision Transformer (ViT) architecture as the image encoder, with a patch size of 16x16 pixels [32].
4.2 Evaluating OnlyFlow
Quantitative assessments
To empirically validate OnlyFlow, we perform extensive testing using a subset of 70,000 unseen videos from the WebVid [1] dataset to generate the optical flow to condition our proposed model. All the assessment was performed on generated videos with $F=16$ frames at a resolution of $512\times 512$ pixels.
As for the evaluation metrics, we validate our model using the FVD, flow fidelity, and textual alignment metrics, previously detailed. Finally, to analyze the impact of our proposed motion module, we evaluate several optical flow conditioning strengths $\gamma$ (Eq. (2)), as well as using two OnlyFlow variants: OnlyFlow-T, referring to injecting the feature in the temporal layers (i.e. our OnlyFlow base model), and OnlyFlow-ST, where we inject the motion features in both spatial and temporal layers.
First, we discuss the FVD metric in Fig. 4(a). Interestingly, we notice that as we increase optical flow conditioning strength $\gamma$, the FVD decreases. We posit that this is explained by the addition of real data as a condition to video generation. Real data brings realistic motion which will end up in being used by OnlyFlow, which in turn will be interpreted as accurate by the FVD metric. In this generative process, enforcing the optical flow of an input video with realistic motion naturally improves the FVD.
@@ 4.2 Evaluating OnlyFlow
Next, we examine the textual alignment results in Fig. 4(c). The outcome of this experiment show that the model produces stable CLIP scores when using any level of optical flow conditioning. This suggests that the content generated by the model is aligned with the text to the same extent, whether the motion cues are used or not.
Finally, we opt to insert the optical flow conditioning representations into the temporal attentional blocks. This decision is motivated by the ability of the temporal attention layer to capture temporal dependencies, which is consistent with the sequential and causal nature of optical flow. From previous empirical results, we find that injecting the motion feature into the spatial and temporal layers together has no advantage over introducing it into the temporal layers, justifying our choice of design injection. For the remainder of the paper, all results were performed using OnlyFlow-T.
@@
Figure 5: Qualitative comparison of video-to-video generation models. Videos are generated using same text prompt and input video. OnlyFlow exhibits a superior combination of motion fidelity and image realism. It positively compares to approaches that use depth maps (RAVE, VideoComposer, Control-A-Video), is comparable to Gen-1 [12]’s temporal coherence and VideoComposer [42]’s image quality.
User preference study
While automatic metrics approximate the behavior of T2V pipelines, the gold standard for evaluating generative models today is a human study. To this end, we use the Two-alternative Forced Choice (2AFC) protocol for text-driven video editing [2, 11, 13, 31, 49]. Participants are presented with the input video, our results, and a baseline, and are asked to decide which video better matches the text prompt, which video better preserves the motion of the input video, and which video they prefer overall. We collected 300 user judgments. As seen in Tab.1, our method is consistently preferred over all baselines.
Qualitative comparison with state-of-the-art video-to-video generation models
For the loosely-defined problem of video-to-video editing and customization, we show that our OnlyFlow model performs better for tasks that require only the motion of the input video. The visual comparison in Fig. 5 reveals that OnlyFlow can generate animations whose movements are close to the input video, given a different textual prompt. In this comparison, the rippling waves effect of the auxiliary video struggle to be accurately represented in other models, except for Control-A-Video and Gen-1. We notice that the prompt adherence is also better followed in our approach, Gen-1, VideoComposer and Tune-A-Video. Nevertheless, the latter two models failed to take motion inspiration from our auxiliary video and Gen-1 seems to stack mountains and ripples visuals. Compared to other work, such as Tune-A-Video [46], our approach is lightweight. On top of that, it does not require any costly DDIM [38] inversion process used in TokenFlow [13] and RAVE [23].
Figure 6: Qualitative comparison on camera-movement controlled video generation. All videos are generated using the same text prompt, camera trajectory (pan-left), and seed when available. Without having trained for this task, our OnlyFlow model achieve the same camera control capability as other camera-movement approaches trained on this task.
In recent work, such as MotionCtrl [44] and CameraCtrl [18], the conditioning of camera motion is achieved with a camera position encoding, while the AnimateDiff [15] motion conditioning uses a LoRA trained on the temporal layer for a given motion. For camera displacements, OnlyFlow provides a flexible alternative for modeling camera movement. To do this, we create an artificial optical flow to describe the desired motion or provide a video with the desired motion. In Fig. 6 we show a specific scenario where we prompted our model with a constant horizontal movement to the left. The generated results clearly show that the generation is on par with the controllability of specific approaches such as CameraCtrl or MotionCtrl. Note that we did not observe satisfactory results with AnimateDiff’s motion LoRA, both in terms of aesthetics and expected camera movements.
4.3 OnlyFlow for Art
Figure 7: One of our interests is the creation of artistic videos. Here we show an example of our results, which we projected in ARTECHOUSE’s art venue.
@@ 4.3 OnlyFlow for Art
As previously shown, OnlyFlow has many appealing properties. Consequently, our model may be applied to many creative scenarios. In particular, we are interested in using OnlyFlow for artistic applications. Using the natural motion present in video footage allows the generation of videos that would require tremendous editing efforts to be created without OnlyFlow. In Fig. 7 and the supplementary video, we provide an artistic video we created during the project, that incorporate upscaling with ESRGAN [41] and frame interpolation with FILM [33], which by themselves showcase the type of outputs that can be done using OnlyFlow. Moreover, OnlyFlow is readily compatible with extensions of AnimateDiff and Stable-Diffusion that are often used in artistic creation context such as IP-Adapter [50], which can further be used to improve the quality and controllability of the generated videos. Finally, we did an artwork based on OnlyFlow over the course of the 2025 SUBMERGE project. It has been presented with ARTECHOUSE’s state-of-the-art 270-degree, 18K-resolution digital canvas in their New York City immersive in-person experience and is included in the CVPR AI Art Gallery 2025.
4.4 Analyzing OnlyFlow
Figure 8: Illustration of OnlyFlow’s optical flow conditioning strength impact on generated videos. Setting the $\gamma$ scale at which the features from the input video are injected sets the influence of its motion on the generated video. The videos presented in each row are obtained for increasing values of $\gamma$ between 0 and $1.0$. We observe a progressive alignment of the generated video motion and wind turbine position on the auxiliary video
Flow strength control
Semantic alignment
We noticed an interesting capability of semantic alignment between the descriptive information from the optical flow fed into the model and the given prompt. As shown in Fig. 9, the optical flow inadvertently captures object shapes. When there is a match between the input video motion concept and the prompt, the model successfully maps the zones of interest to be generated. This ability is similar to other conditioning models such as ControlNet [52] and T2I-Adapter [30], but our method does not require manual control map creation.
(a) Frames from the auxiliary video of which optical flow is extracted
(b) Frames from corresponding generated video
Figure 9: Example of semantic alignment between optical flow and prompt. With the prompt “Trees in forest”, our OnlyFlow model generates a clothesline corresponding to the teeth of the smile.
5 Conclusions
In this paper, we introduced our OnlyFlow method, which improves on T2V models such as AnimateDiff by conditioning it on the motion extracted from an input auxiliary video. The optical flow of the auxiliary video is then passed to a trainable flow encoder connected to the T2V model. The flow encoder outputs features maps which are injected into the attention layers of the U-net. The generated video is encouraged to accurately follow the motion of the input video. While other methods allow conditioning by movement (with motion masks or user-defined vectorial camera movements), OnlyFlow allows generating any video with a text prompt and the motion of an auxiliary video. In our experiments, we demonstrate the efficiency of OnlyFlow for various applications such as camera motion control (panning, zooming, etc.) and video editing/style transfer. Without necessarily being trained for such applications, OnlyFlow compares positively to other state-of-the-art methods that have been specifically trained for such applications. We also carried out a user study which confirmed the interest of our method in generating very interesting videos, demonstrating that their movement convincingly and coherently follows that of the input reference.
@@ 5 Conclusions
One clear limitation of our work is photorealism obtained in the generated videos. While T2V models like AnimateDiff [15] are powerful generative models, their generation remains poor for most complex prompts depicting scenes with various objects and subjects or scenes with movements that might not appear in the training set. The resolution of the generated video is also limited by the underlying model.
@@
6 Acknowledgments
This project was provided with computing HPC & AI and storage resources by GENCI at IDRIS thanks to the grant 2024-AD011014329R1 on the supercomputer Jean Zay’s V100 & A100 partitions.
This research was funded by the French National Research Agency (ANR) under the project ANR-23-CE23-0023 as part of the France 2030 initiative.
Supplementary Material
7 Implementation Details
Dataset
We use a random horizontal flip for videos with a 50 percent probability and randomly crop a 256 by 384 area out of the spatially downsampled files.
Optimizer.
We used the Adam optimizer [25] with a constant learning rate of $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=5e-8$ for numerical stability and a weight decay of $10^{-2}$. For each parameter update, we clip the gradient norm to $0.4$.
Optical flow Feature Extraction.
The optical flow encoder first unshuffles the input video by a ratio of $8$, increasing the number of channels. For each of the output channels resolution $(320,640,1280,1280)$ we want to obtain, the encoder proceeds in a cascading way, applying $2$ times the following blocks:
  • a ResNet block with a downsampling layer
  • a temporal attention block containing $8$ heads, with a sinusoidal positional embedding on $16$ frames
Sampling.
We use a PNDM scheduler [29] with a linear beta schedule where $\beta_{start}=0.00085$, $\beta_{end}=0.012$, and $T=1000$. To allow classifier-free guidance, we randomly drop the text condition 10% of the time.
RAFT settings.
In both training and evaluation phases we used the RAFT large checkpoint with the defaults number of $12$ optical flow refining iteration updates.
8 Usage tips and tricks
Input video frame rate
Our model inference contains two opposing constraints on the conditioning frame rate. On one side, optical flow estimation model perform better between frames that are similar, meaning higher frame rate. At the same time, T2V models often generate 16 or 24 frames per forward pass. Training datasets like WebVid often correspond to video downsampled temporally to 8 fps.
If the optical flow given in input to our model is not within the range of motion of what the base T2V model can achieve in its generation, we may observe a deterioration in prompt adherence or realism.
Aspect ratio.
As the AnimateDiff model allows it, we can generate non squared video. Nevertheless, because of the non-convolutional nature of our flow encoder, the optical flow dimensions have to be a multiple of 64.
9 User study details
We submitted the following question to the panel of participants expressing their choices for each pair of video :
  • Which video best respect the input text?
  • Which video best replicate the motion from the input video?
  • Which video do you prefer overall?